Fermi National Laboratory
 

Operation Summary for October 14, 2011 - October 17, 2011

 

NuMI reported receiving 9.1E18 protons on target for the period from 10/10/11 to 10/17/11.

Friday October 14

The day shift began with Operations monitoring the protons in the Recycler (4.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Neutron Therapy Facility (NTF) preparing to treat a patient later in the morning, with the MiniBooNE experiment off in preparation for an MI-12 access, and with the NuMI experiment taking beam.

During the day shift, Operations established beam to the Meson FTBF experiment T-992 at 10:40 AM. MiniBooNE came out of access and Operations established beam to the experiment at 10:49 AM. A Booster person began a study at 11:01 AM. NTF personnel reported at 1:54 PM that they had completed their patient treatment. At 2:01 PM, Power Supply personnel began taking the MI-40 lower power supply out of the circuit for work; this work held off beam to MiniBooNE and T-992 for about 12 minutes. At the same time as this work, Booster RF personnel began work on a RF station (BRF12); they brought it back online about 18 minutes later. At 2:15 PM, two Linac RF stations (LRF4 & 5) tripped off. Linac personnel investigated and reported that LRF5 had blown a fuse in its modulator; they had the fuse replaced by 3:22 PM. An Antiproton Source (Pbar) person began a study at 3:41 PM. MTA personnel began a study at the end of the shift.

During the evening shift, MTA personnel completed their studies at 5:08 PM. A Pbar person completed his study at 7:24 PM. Operations disabled beam to T-992 around 9:15 PM.

Saturday October 15

The midnight shift began with Operations monitoring protons in the Recycler (10.5E10) and by sending beam to NuMI and MiniBooNE.

During the midnight shift, at 4:43 AM, a MI Lambertson magnet power supply tripped off on excessive ripple; it reset. Operators reported short downtimes due to BRF7 trips.

During the day shift, at 8:59 AM, a NuMI magnet (E:HV101) tripped off on a high temperature indication for the first of two times in 14 minutes; it reset. Operators contacted a NuMI person. Operations turned off beam to MI at 10:06 AM and operators prepared MI for an access to check the magnet’s temperature. EE Support personnel will use this time to work on a MI-30 VCB. Operators came out of the MI access at 12:01 PM and reported that it was difficult to get good temperature readings on HV101. Operations had beam established to MI, NuMI, and MiniBooNE by 12:59 PM. At 3:26 PM, HV101 tripped off on a high temperature indication; it reset. At 3:45 PM, a MI Lambertson magnet (I:LAM42) power supply tripped off on excessive ripple; it reset.

During the evening shift, at 4:12 PM, I:LAM42 tripped off on excessive ripple; it reset. Operations established beam to T-992 at 4:15 PM. At 9:56 PM, operators reported that the I- Source output current was dropping; operators tried to recover the source and called in Linac personnel for help. Operations disabled beam to T-992 at 10:02 PM.

Sunday October 16

The midnight shift began with Operations monitoring protons in the Recycler and sending beam to NuMI and MiniBooNE.

During the midnight shift, a NuMI cooling system person reported at 12:23 AM that he had found a correlation between the HV101 magnet temperature trips and the NuMI target hall water skid room temperature; others pointed out that this might only be a coincidence; the investigation continued. Linac personnel’s tuning on the I- Source had its output back up to normal by 12:28 AM. I:LAM42 tripped off on excessive ripple at 12:56 AM; it reset. Operations established beam to T-992 at 6:09 AM.

During the day shift, at 10:58 AM, I:LAM42 tripped off on excessive ripple; it reset. Operators tuned on the Booster and MI machines all shift.

During the evening shift, Operations disabled beam to T-992 at 9:03 PM. At 11:05 PM, BRF7 tripped off for the first of three times in 52 minutes; it reset. At 12:59 PM, I:LAM42 tripped off on excessive ripple; it reset.

Monday October 17

The midnight shift began with Operations monitoring protons in the Recycler and sending beam to NuMI and MiniBooNE.

During the midnight shift, at 2:17 AM, BRF7 tripped off for the first of eight times in about two and a half hours; it reset. Operators lowered BRF7’s modulator anode voltage. At 6:49 AM, I:LAM42 tripped off on excessive ripple; it reset.

Operations Coordinator (OPSCO)

The OpsCo asked MI, NuMI, and mechanical personnel to continue the investigation of the HV101 temperature trips.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/17/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory